# Yintong Huo‡§, Michael R. Lyu

### v1

@@ Abstract
Index Terms:
Multimodal Large Language Models, Code Generation, Web Development.
I Introduction
Converting webpage designs into functional UI code is a critical yet labor-intensive step in web development. MLLMs have demonstrated remarkable performance on visually rich code generation tasks [1, 2, 3, 4, 5], creating new opportunities for automated design-to-code conversion that replicates webpage elements, layouts, text, and colors.
Several benchmarks have been proposed to evaluate MLLMs on front-end code generation. These benchmarks typically involve either synthesizing webpage code via LLMs (e.g., WebSight [6] and Web2Code [7]), or curating code by cleaning real-world webpages (e.g., Webcode2M [8] and Design2Code [9]). However, while these evaluation benchmarks are effective in measuring certain MLLM capabilities, they do not adequately represent the complex challenges faced by developers in real-world development scenarios.
(1) Lack of front-end framework integration. Current benchmarks fail to incorporate front-end frameworks, such as React, Vue, and Angular, which are integral to modern web development (as indicated in 2025 trends [10]). Consequently, MLLM capabilities for practical framework-based website development remain unexplored.
(2) Insufficient task coverage. Existing benchmarks inadequately cover the comprehensive spectrum of front-end development tasks. As illustrated in Fig. 1, development encompasses not only initial code generation from UI designs, but also iterative code editing for design refinement and repairing design-related issues [11, 12]. However, current evaluations focus exclusively on the generation phase, neglecting the critical edit and repair phases.
Figure 1: Pipeline of MLLM-based automated front-end engineering.
@@ I Introduction
TABLE I: Comparison of existing UI-to-Code benchmark and DesignBench. Vanilla refers to plain HTML/CSS.
@@
Pix2code [13] 1742 Synthetic Vanilla Generation
WebSight [6] 823K Synthetic Vanilla Generation
Web2Code [14] 60K Synthetic Vanilla Generation
WebCode2M [8] 20K Real-world Vanilla Generation
Design2Code [9] 484 Real-world Vanilla Generation
DesignBench (Ours) 900 Real-world Vanilla, React, Vue, Angular Generation, Edit, Repair Difficulty, Context, Code
@@
  • We introduce the first comprehensive multi-framework, multi-task benchmark for evaluating MLLMs’ capabilities in automated front-end engineering across HTML/CSS, React, Vue, and Angular frameworks.
@@
  • We reveal key insights into MLLMs’ framework-specific limitations, identify task-dependent performance bottlenecks, and characterize performance variations under diverse conditions. These results provide important guidance for future research and practice in front-end development.
  • To facilitate further research, we publicly release the code and data on Github [15].
II Background
II-A Web Development Process
A typical web application front-end development process includes the following stages:
  1. 1.
    Design stage: designers create high-fidelity mock-ups using prototyping tools such as Sketch [16] and Axure [17] during design stage.
  2. 2.
@@ II-A Web Development Process
    Development stage: this phase involves transforming the design concepts into a functional application through coding. The development stage typically consists of the implementation of the GUI and the underlying functionalities. As shown in Fig. 1, the developer first prompts the MLLM to generate the front-end code based on the UI-Mockup. However, the version one code did not fully comply with the UI-Mockup. Then the user asked MLLM to edit the code to change the background color and add buttons to generate version two. Then during the actual deployment process, the front-end engineer found a display issue on the UI, that is, a collision between the avatar and the name, and asked MLLM to fix it and generate version three code.
II-B Front-end Web Development Framework
@@
Front-end frameworks are collections of pre-written, reusable code that provide a foundation for building the user interface of a website. These frameworks can save time and improve the quality of the final product by providing a standardized set of components and templates for front-end development. In 2025, React, Vue, and Angular have emerged as the predominant frameworks in web development [10], with 39.5%, 15.4%, and 17.1% popularity [18], respectively.
III Related Work
III-A UI Code Generation
@@
UI code generation techniques fall into three categories: Deep Learning (DL) based methods, Computer Vision (CV) based methods, and Multimodal Large Language Model (MLLM) based methods. (1) DL-based methods: [19, 20, 21, 22, 23] leverages CNNs to automatically prototype software GUIs. Pix2code [13] utilizes CNNs and LSTM to extract features from GUI images to generate a domain-specific language (DSL). [24] implements an encoder-decoder framework with an attention mechanism to generate the DSL. (2) CV-based methods: Sketch2Code [25] inputs hand-drawn sketches into object detection models to learn the object representation, which is read by the UI parser to generate code for targeted platforms. REMAUI [26] identifies user interface elements via optical character recognition (OCR) techniques and then infers a suitable user interface hierarchy and exports the results as source code. (3) MLLM-based methods [27, 28, 29, 30]: to solve the element omission distortion and misarrangement problems during UI code generation, DCGen [28] proposes a divide-and-conquer-based approach to generate the code of the submodules separately and then assemble them to construct the full webpage based on MLLMs. DeclarUI [29] applies the element segmentation method to accurately generate elements and page transition graphs to prompt MLLMs to generate mobile app UI with jump logic. While prior works achieve decent performance in UI-to-code, none of them address the conventional framework-based development and other essential front-end procedures like design edit and repair.
III-B Benchmarks for UI Code Generation
@@
Many benchmarks [31, 32, 33, 34] have been proposed to evaluate the code generation, understanding and reasoning capabilities of LLMs, but there are only a limited number of benchmarks for UI code. To improve MLLMs’ UI-to-Code capabilities, Pix2code introduced a domain-specific language (DSL) and leveraged CNN and LSTM architectures to translate UI mockups into DSL code. However, due to the inherent limitations of DSL approaches, Pix2code suffers from poor scalability and limited real-world applicability. WebSight advanced the field by synthesizing high-quality HTML code for training, while Web2Code [7] proposed the Webpage Code Generation Benchmark (WCGB) to systematically evaluate MLLMs’ HTML parsing capabilities. Despite these contributions, both benchmarks rely on synthetic data, which may not capture the complexity and variability of real-world web development scenarios. Design2Code [9] addressed this limitation by manually curating 484 authentic web pages from the Common Crawl dataset, constructing the first real-world benchmark for design-to-code evaluation. Building upon this foundation, WebCode2M [8] significantly expanded the scale with 20,000 samples, providing both comprehensive training data for model development and robust test sets for evaluation. Other Benchmarks like Interaction2Code [35] and MRWeb [36] mainly focus on the interactive and multi-page web application generation, which is out of our scope.
@@
Table I presents a comparison between existing benchmarks and our, in terms of the size, collection method, framework, target tasks, and evaluation dimension. DesignBench is distinct in incorporating varied front-end frameworks, multiple tasks, and diverse evaluation dimensions.
IV DesignBench
IV-A Task Definition
@@
IV-B Data Collection
Figure 2: Pipeline of DesignBench  construction.
Design Generation. For websites developed by frameworks, we collect samples from GitHub and the top 500 globally visited websites. Existing reports [10, 18] show that the three most popular front-end development frameworks are React, Vue, and Angular, so we collect data for these three frameworks to build our benchmark. (1) Github projects. We search for “React projects”, “Vue projects” and “Angular projects” to get a summary list of web projects with different frameworks, then we identify 152 popular projects with deployed links and higher star counts. These projects represent various real-world website uses, ranging from commercial product front-end websites to blogs, with 4055 average star counts and an average of 996 forks. Their popularity has proven their usefulness and quality. (2) Top 500 globally visited websites. These websites are ranked by Moz [37], and we collect 158 webpages that adopt frameworks as our dataset.
For websites without applying frameworks, we sample 120 webpages with different lengths from the webcode2m dataset [8]. Ultimately, we compile a dataset consisting of 430 webpages. After getting the link of the webpage, we apply single-file tool [38] to save the webpage as a stand-alone file, then replace the images in the web page with placeholders, and finally use selenium [39] to take screenshots for constructing the input design image $I$ for task $\mathcal{T}_{G}$.
Design Edit. To obtain real-world user instructions for UI design modification, we crawl 541 React-based projects from Vercel’s V0 platform [40] and 1,349 Vue-based projects from Vue0 [41]. These projects contain comprehensive interaction histories, including user instructions and corresponding UI code implementations.
@@ IV-B Data Collection
However, some edit steps feature ambiguous user instructions or poorly UI modifications that fail to satisfy the intended requirements. To filter out low-quality samples, we employ five PhD students with three years of front-end development experience to conduct a comprehensive assessment.
Specifically, the annotators classify user instructions into three levels of clarity: clear, moderate, and ambiguous, based on detailed assessment of instruction specificity, actionability, and comprehensibility. Similarly, they evaluate the quality of modified UI and categorize them into terrible, good, and excellent based on how well the changes align with the given instructions. The detailed annotation guidelines are in [15]. The final classifications for both instruction clarity and UI modification effectiveness are determined through a majority voting process among the annotators.
Finally, we curate 359 high-quality React and Vue samples that receive clear instruction ratings and excellent modification quality scores. To obtain Angular and vanilla HTML/CSS samples, we randomly select 146 samples from this curated set and translate them into Angular and vanilla HTML/CSS code implementations using GPT-4o [42]. Following the automated translation, the same five PhD students verify the correctness of the translated implementations, and make necessary modifications to ensure the code meet users’ instructions.
Design Repair. After collecting the webpages of the above two tasks, the five PhD students screen out the UI samples with design issues. They follow comprehensive guidelines to identify display problems, including: (1) Layout issues such as misaligned elements and incorrect positioning; (2) Visual inconsistencies including improper spacing, incorrect font sizes, or color scheme violations; (3) Component rendering problems such as missing elements, overlapping content, or distorted images. The detailed annotation guidelines are in [15].
Finally, we get 111 webpages with problematic UI code $C_{p}$ and image $I_{p}$. To obtain the repaired code $C_{r}$ for evaluation, 111 webpages are evenly assigned to the five PhD students to manually fix the UI display issues. Each student spent about 5 hours to complete the repair of all assigned samples.
IV-C Data Annotation
Figure 3: The UI issue types in DesignBench. The red bounding box marks the issue location.
@@ IV-C Data Annotation
Design Repair. The annotators are employed to annotate the UI issues. Aftering collecting the webpages with display issues, we first randomly select 25% samples for analysis and then discuss, revise, and refine the UI issue type until everyone reaches a consensus. During annotating new data, if encountering a new issue type, annotators will communicate and update issue type in time to guide subsequent annotations. Finally, we classify 6 types of UI display issues as follows:
@@
IV-D Benchmark Statistics
Table II presents the sample counts of the three tasks with three frameworks, the Vanilla denotes the webpage developed by vanilla HTML/CSS.
TABLE II: Sample counts of three frameworks on three tasks.
@@ IV-D Benchmark Statistics
(b) Issue Distribution.
Figure 4: Category distribution of Design Generation and Design Repair task.
(a) Operation Distribution.
(b) Visual Distribution.
Figure 5: Edit type distribution of Design Edit task.
@@
Fig. 4(b) shows the distribution of the UI issues of design repair task. The issue types span multiple dimensions with alignment being most prevalent (42.2%), followed by crowding (18.7%), occlusion (18.1%), overflow (11.4%), and other visual defects. This wide range of issue types demonstrates that design repair tasks address comprehensive quality assurance across multiple design issues dimensions.
V Experiment Setup
V-A Models
@@
The studied MLLMs are listed in Table III. We select six state-of-the-art LLMs that have been widely explored in multimodal tasks, three from open-source models, namely Pixtral [43], Qwen [44], and LLama [45], three from commercial models like Gemini [46], GPT [42] and Claude [47]. In configuring the MLLM models, we set the temperature to 0 and the maximum number of tokens output as the upper limit of MLLMs’ maximum output token. All other parameters are kept at their default settings. The detailed prompts of the three tasks are available in our code repository [15]. The entire benchmark evaluation incurs an average API cost of \$52 per model on average. The average processing times per sample with a single thread, are 49 seconds for generation, 29 seconds for editing, and 25 seconds for repair, respectively.
V-B Metric
We evaluate the performance of the model on DesignBench from three types of metrics:

### v2

# DesignBench: A Comprehensive Benchmark for MLLM-based Front-end Code Generation

@@ Abstract.
Multi-modal Large Language Model, Code Generation, Web Development
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: July, 2026; DC, USA; isbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Automatic programmingccs: Computing methodologies Artificial intelligence
1. Introduction
Converting webpage designs into functional UI code is a critical yet labor-intensive step in web development. MLLMs have demonstrated remarkable performance on visually rich code generation tasks (Yang et al., 2025; Liu et al., 2025; Zhang et al., 2025; Li et al., 2024; Chen et al., 2020; Tang et al., 2025), creating new opportunities for automated design-to-code conversion that replicates webpage elements, layouts, text, and colors.
Several benchmarks have been proposed to evaluate MLLMs on front-end code generation. These benchmarks typically involve either synthesizing webpage code via LLMs (e.g., WebSight (laurençon2024unlocking) and Web2Code (Yun et al., 2024)), or curating code by cleaning real-world webpages (e.g., Webcode2M (Gui et al., 2025) and Design2Code (Si et al., 2024)). However, while these evaluation benchmarks are effective in measuring certain MLLM capabilities, they do not adequately represent the complex challenges faced by developers in real-world development scenarios.
(1) Lack of front-end framework integration. Current benchmarks fail to incorporate front-end frameworks, such as React, Vue, and Angular, which are integral to modern web development (as indicated in 2025 trends (Raju, 2025)). Consequently, MLLMs’ capabilities for practical framework-based website development remain unexplored.
(2) Insufficient task coverage. Existing benchmarks inadequately cover the comprehensive spectrum of front-end development tasks. As illustrated in Fig. 1, development encompasses not only initial code generation from UI designs, but also iterative code editing for design refinement and repairing design-related issues (Liu et al., 2020, 2022). However, current evaluations focus exclusively on the generation phase, neglecting the critical edit and repair phases.
Figure 1. Pipeline of MLLM-based automated front-end engineering.
@@ 1. Introduction
Table 1. Comparison of existing UI-to-Code benchmark and DesignBench. Vanilla refers to plain HTML/CSS.
@@
Pix2code (Beltramelli, 2018) 1742 Synthetic Vanilla Generation
WebSight (laurençon2024unlocking) 823K Synthetic Vanilla Generation
Web2Code (Gui et al., 2024) 60K Synthetic Vanilla Generation
WebCode2M (Gui et al., 2025) 20K Real-world Vanilla Generation
Design2Code (Si et al., 2024) 484 Real-world Vanilla Generation
DesignBench 900 Real-world Vanilla/React/Vue/Angular Generation/Edit/Repair Difficulty/Context/Code/Failure Type
@@
  • We introduce the first multi-framework, multi-task benchmark for evaluating MLLMs’ capabilities in automated front-end engineering across HTML/CSS, React, Vue, and Angular frameworks.
@@
  • We reveal key insights into MLLMs’ framework-specific limitations, identify task-dependent performance bottlenecks, characterize performance variations under diverse conditions and identify 22 types of failure of the three tasks. These results provide important guidance for future research and practice in front-end development.
2. Background
2.1. Web Development Process
@@
Web application front-end development process includes the following stages: (1) Design stage: designers create high-fidelity mock-ups using prototyping tools such as Sketch (37) and Axure (5) during design stage. (2) Development stage: it involves transforming the design concepts into a functional application through coding. The development stage typically consists of the implementation of the GUI and the underlying functionalities. As shown in Fig. 1, the developer first prompts the MLLM to generate the front-end code based on the UI-Mockup. However, the version one code did not fully comply with the UI-Mockup. Then the user asked MLLM to edit the code to change the background color and add buttons to generate version two. Then during the actual deployment process, the front-end engineer found a display issue on the UI, that is, a collision between the avatar and the name, and asked MLLM to fix it and generate version three code.
2.2. Front-end Web Development Framework
@@
Front-end frameworks are collections of pre-written, reusable code that provide a foundation for building the user interface of a website. These frameworks can save time and improve the quality of the final product by providing a standardized set of components and templates for front-end development. In 2025, React, Vue, and Angular have emerged as the predominant frameworks in web development (Raju, 2025), with 39.5%, 15.4%, and 17.1% popularity (Krotoff, 2024), respectively.
3. Related Work
3.1. UI Code Generation
@@
UI code generation techniques fall into three categories: Deep Learning (DL) based methods, Computer Vision (CV) based methods, and Multimodal Large Language Model (MLLM) based methods. (1) DL-based methods: (Aşıroğlu et al., 2019; Cizotto et al., 2023; Moran et al., 2018; Xu et al., 2021; Chen et al., 2018) leverages CNNs to automatically prototype software GUIs. Pix2code (Beltramelli, 2018) utilizes CNNs and LSTM to extract features from GUI images to generate a domain-specific language (DSL). (Chen et al., 2022) implements an encoder-decoder framework with an attention mechanism to generate the DSL. (2) CV-based methods: Sketch2Code (Jain et al., 2019) inputs hand-drawn sketches into object detection models to learn the object representation, which is read by the UI parser to generate code for targeted platforms. REMAUI (Nguyen and Csallner, 2015) identifies user interface elements via optical character recognition (OCR) techniques and then infers a suitable user interface hierarchy and exports the results as source code. (3) MLLM-based methods (Lu et al., 2025; Wan et al., 2025b; Zhou et al., 2025; Yuan et al., 2025; Wan et al., 2025a; Xiao et al., 2025, 2026; Dang et al., 2025): to solve the element omission distortion and misarrangement problems during UI code generation, DCGen (Wan et al., 2025b) proposes a divide-and-conquer-based approach to generate the code of the submodules separately and then assemble them to construct the full webpage based on MLLMs. DeclarUI (Zhou et al., 2025) applies the element segmentation method to accurately generate elements and page transition graphs to prompt MLLMs to generate mobile app UI with jump logic. EfficientUICoder (Xiao et al., 2025) accelerates UI code generation by input and output token compression. ComUICoder (Xiao et al., 2026) apply semantic-aware segmentation and merging methods for reusable UI code generation. While prior works achieve decent performance in UI-to-code, none of them address the conventional framework-based development and other essential front-end procedures like design edit and repair.
3.2. Benchmarks for UI Code Generation
@@
Many benchmarks (Gu et al., 2024; Lam et al., 2025; Jain et al., 2024; Zhuo et al., 2024) have been proposed to evaluate the code generation, understanding and reasoning capabilities of LLMs, but there are only a limited number of benchmarks for UI code. To improve MLLMs’ UI-to-Code capabilities, Pix2code introduced a domain-specific language (DSL) and leveraged CNN and LSTM architectures to translate UI mockups into DSL code. However, due to the inherent limitations of DSL approaches, Pix2code suffers from poor scalability and limited real-world applicability. WebSight advanced the field by synthesizing high-quality HTML code for training, while Web2Code (Yun et al., 2024) proposed the Webpage Code Generation Benchmark (WCGB) to systematically evaluate MLLMs’ HTML parsing capabilities. Despite these contributions, both benchmarks rely on synthetic data, which may not capture the complexity and variability of real-world web development scenarios. Design2Code (Si et al., 2024) addressed this limitation by manually curating 484 authentic web pages from the Common Crawl dataset, constructing the first real-world benchmark for design-to-code evaluation. Building upon this foundation, WebCode2M (Gui et al., 2025) significantly expanded the scale with 20,000 samples, providing both comprehensive training data for model development and robust test sets for evaluation. Other Benchmarks like Interaction2Code (Xiao et al., 2024) and MRWeb (Wan et al., 2024) mainly focus on the interactive and multi-page web application generation, which is out of our scope.
@@
Table 1 presents a comparison between existing benchmarks and our, in terms of the size, collection method, framework, target tasks, and evaluation dimension. DesignBench is distinct in incorporating varied front-end frameworks, multiple tasks, and diverse evaluation dimensions.
4. DesignBench
4.1. Task Definition
@@
4.2. Data Collection
Figure 2. Pipeline of DesignBench construction.
Design Generation. For websites developed by frameworks, we collect samples from GitHub and the top 500 globally visited websites. Existing reports (Raju, 2025; Krotoff, 2024) show that the three most popular front-end development frameworks are React, Vue, and Angular, so we collect data for these three frameworks to build our benchmark. (1) Github projects. We search for “React projects”, “Vue projects” and “Angular projects” to get a summary list of web projects with different frameworks, then we identify 152 popular projects with deployed links and higher star counts. These projects represent various real-world website uses, ranging from commercial product front-end websites to blogs, with 4055 average star counts and an average of 996 forks. Their popularity has proven their usefulness and quality. (2) Top 500 globally visited websites. These websites are ranked by Moz (41), and we collect 158 webpages that adopt frameworks as our dataset.
For websites without applying frameworks, we sample 120 webpages with different lengths from the webcode2m dataset (Gui et al., 2025). Ultimately, we compile a dataset consisting of 430 webpages. After getting the link of the webpage, we apply single-file tool (Lormeau, 2023) to save the webpage as a stand-alone file, then replace the images in the web page with placeholders, and finally use selenium (Selenium, 2025) to take screenshots for constructing the input design image $I$ for task $\mathcal{T}_{G}$.
Design Edit. To obtain real-world user instructions for UI design modification, we crawl 541 React-based projects from Vercel’s V0 platform (42) and 1,349 Vue-based projects from Vue0 (43). These projects contain comprehensive interaction histories, including user instructions and corresponding UI code implementations.
@@ 4.2. Data Collection
However, some edit steps feature ambiguous user instructions or poorly UI modifications that fail to satisfy the intended requirements (here [^1] https://github.com/WebPAI/DesignBench/blob/main/assets/Unclear_Failed_Examples.md are some bad cases). To filter out low-quality samples, we employ five PhD students with three years of front-end development experience to conduct a comprehensive assessment.
Specifically, the annotators classify user instructions into three levels of clarity: clear, moderate, and ambiguous, based on detailed assessment of instruction specificity, actionability, and comprehensibility. Similarly, they evaluate the quality of modified UI and categorize them into terrible, good, and excellent based on how well the changes align with the given instructions. The detailed annotation guidelines are in (Team, ). The final classifications for both instruction clarity and UI modification effectiveness are determined through a majority voting process among the annotators.
Finally, we curate 359 high-quality React and Vue samples that receive clear instruction ratings and excellent modification quality scores. To obtain Angular and vanilla HTML/CSS samples, we randomly select 146 samples from this curated set and translate them into Angular and vanilla HTML/CSS code implementations using GPT-4o (OpenAI, 2024). Following the automated translation, the same five PhD students verify the correctness of the translated implementations, and make necessary modifications to ensure the code meet users’ instructions.
Design Repair. After collecting the webpages of the above two tasks, the five PhD students filter the samples with design issues. They follow comprehensive guidelines to identify display issues, including: (1) Layout issues such as misaligned elements and incorrect positioning; (2) Visual inconsistencies including improper spacing, incorrect font sizes, or color scheme violations; (3) Component rendering problems such as missing elements, overlapping content, or distorted images.
Finally, we get 111 webpages with problematic UI code $C_{p}$ and image $I_{p}$. To obtain the repaired code $C_{r}$ for evaluation, 111 webpages are evenly assigned to the five PhD students to manually fix the issues. Each student spent about 5 hours to complete the repair of all assigned samples. The detailed annotation guidelines of these three tasks are in our repo [^2] https://github.com/WebPAI/DesignBench/blob/main/assets/Guideline.md.
4.3. Data Annotation
Figure 3. The UI issue types in DesignBench. The red bounding box marks the issue location.
@@ 4.3. Data Annotation
Design Repair. The annotators are employed to annotate the UI issues. After collecting the webpages with display issues, we first randomly select 25% samples for analysis and then discuss, revise, and refine the UI issue type until everyone reaches a consensus. During annotating new data, if encountering a new issue type, annotators will communicate and update issue type in time to guide subsequent annotations. Finally, we classify 6 types of UI display issues as follows:
@@
4.4. Benchmark Statistics
Table 2 presents the sample counts of the three tasks with three frameworks, the Vanilla denotes the webpage developed by vanilla HTML/CSS.
Table 2. Sample counts of three frameworks on three tasks.
@@ 4.4. Benchmark Statistics
(b) Operation Distribution.
(c) Visual Distribution.
(d) Issue Distribution.
Figure 4. Type distribution of three tasks.
@@
Fig. 4(d) shows the distribution of the UI issues. The issue types span multiple dimensions with alignment being most prevalent (42.2%), followed by crowding (18.7%), occlusion (18.1%), overflow (11.4%), and other visual defects. This wide range of issue types demonstrates that design repair tasks address comprehensive quality assurance across multiple design issues dimensions.
5. Experiment Setup
5.1. Prompt Design
The simplified prompts for the three tasks are illustrated in Fig.5, with detailed prompts available in our code (Team, ). Our prompts are designed with the following structure: (1) background description: a concise scene setting statement (“You are a senior vue/react/angular developer”, ”You are tasked with generating/editting/repairing”) primes the model for the correct role and task description; (2) detailed and specific task requirements: For design generation, exactly reproduce the screenshot: fonts, colors, layout, spacing. For design edit, surgically alter only the elements specified in the natural-language instruction while leaving all other visuals untouched. For design repair, identify and correct the listed UI defects (occlusion, crowding, overlap, alignment, color/contrast, overflow). (3) framework-specific coding conventions: the prompts guide the model to import available UI components and follow framework-specific coding conventions, with a generic example to reduce ambiguity; (4) output format instructions and examples, each prompt includes strict output format requirements to ensure that the generated code can be reliably extracted and evaluated.
Figure 5. Prompts of design generation, edit and repair task.
5.2. Models
@@
The studied MLLMs are listed in Table 4. We select six state-of-the-art LLMs that have been widely explored in multimodal tasks, three from open-source models, namely Pixtral (Agrawal et al., 2024), Qwen (Bai et al., 2025), and LLama (Meta, 2024), three from commercial models like Gemini (Google, 2024), GPT (OpenAI, 2024) and Claude (Anthropic, 2024). In configuring the MLLM models, we set the temperature to 0 and the maximum number of tokens output as the upper limit of MLLMs’ maximum output token. All other parameters are kept at their default settings. The detailed prompts of the three tasks are available in our code repository (Team, ). The entire benchmark evaluation incurs an average API cost of \$52 per model on average. The average processing times per sample with a single thread, are 49 seconds for generation, 29 seconds for editing, and 25 seconds for repair, respectively.
5.3. Data Contamination Check
First, in terms of the data source, we review the technical reports and training datasets of open source MLLMs. The training datasets for Qwen (Bai et al., 2025), Pixtral (Agrawal et al., 2024) and LLama (Meta, 2024) includes image captions, interleaved image-text data, optical character recognition (OCR) data, visual knowledge, crawling plain text and web document for Q&A, but not web html code. Furthermore, data leakage is particularly unlikely for design edit and design repair, because for design edit task, the training datasets don’t contain “¡instruction, image, react/vue/angular/html code¿” triple corpora. For design repair, there is no ground-truth repair labels are publicly available, as the benchmark is generated and verified by our annotators. To confirm that MLLMs are not “cheating” by memorizing and reproducing the original webpage code, we measure the BLEU score (Papineni et al., 2002) between the generated and original webpage code in the following table. The low BLEU score indicates that the model generates code from the design images rather than directly copying the original webpage.
Table 3. The BLEU score between original webpage code and generated code.
0.1531 0.1318 0.1200 0.0771 0.0562 0.1117 0.1151 0.1338 0.0907
5.4. Metric